OpenStudy (mrs.ambrose614):

The amount of a radioactive material changes with time. The table below shows the amount of radioactive material f(t) left after time t: t(hours) 0 1 2 f(t) 180 90 45 Which exponential function best represents the relationship between f(t) and t? f(t) = 0.5(180)t f(t) = 180(0.25)t f(t) = 180(0.5)t f(t) = 0.5(50)t @confluxepic HELP WILLMEDAL

OpenStudy (irishboy123):

This halves in 1 hour So : \(N = N_o \left( \dfrac{1}{2}\right)^t \implies N = 180 \left( \dfrac{1}{2}\right)^t \)

OpenStudy (mrs.ambrose614):

So is the ANWSER b?????

OpenStudy (sooobored):

\[\frac{1}{2} = 0.5 \ne 0.25\]

OpenStudy (mrs.ambrose614):

So which ANWSER is it??????

OpenStudy (noname):

u only want answer?

OpenStudy (noname):

hello? do you have answer?